A 30 kg child is the only rider of a Ferris wheel with 12 cars and a radius of 10 meters. Each of the carriages of the ride has a mass of 80 kg.



a. What torque does the child exert on the wheel at the 9 o’clock position?


b. the 10 o’clock position?


c. the 11 o’clock position?


d. the 12 o’clock position?

1
Expert's answer
2021-11-26T10:30:08-0500

a)


"\tau=(30)(9.8)(10)\sin{90}=2940\ Nm"

b)


"\tau=(30)(9.8)(10)\sin{60}=2546\ Nm"

c)


"\tau=(30)(9.8)(10)\sin{30}=1470\ Nm"

d)


"\tau=(30)(9.8)(10)\sin{0}=0\ Nm"